3x-2y+5z=31
x+3y-3z=-12
-2x-5y+3z=11

Answer by Edwin McCravy(20064)

3x-2y+5z=31
x+3y-3z=-12
-2x-5y+3z=11
There are bunches of different methods that you study to 
solve systems of equations.  We aren't going to solve them 
every possible way here, and we aren't going to pick a way 
arbitrarily hoping that's the one method you're studying.
So since you didn't tell us which of the many different
methods there are to solve them that you're presently studying,
I'll just give you the solution, which is

(x,y,z) = (3,-1,4)

If you want to know how to solve it a particular way, then post 
again telling us which method you want us to use and we'll show
you how to do it that way.  

Edwin